Please help me solve this equation: 4x^2-6x -4 = 0.
Answer by checkley77(12844)

4x^2-6x -4 = 0
(4x+2)(x-2)=0
4x+2=0
4x=-2
x=-2/4
x=-1/2 ans.
x-2=0
x=2 ans.